A market has a demand function given by the equation Qd= 180- 2p and a supply function given by the equation Qs= -15 + p. The market is government-regulated with price support per unit and production quotas.


a. If the price is set at $72 per unit, what production quota is needed to make sure there are no shortages or surpluses.



1
Expert's answer
2021-10-18T11:28:35-0400

Solution:

First, derive the current market equilibrium price and quantity:

At equilibrium: Qd = Qs

180 – 2P = -15 + P

180 + 15 = P + 2P

195 = 3P

P = 65

Equilibrium price = 65

Substitute in the demand or supply function to derive quantity:

Qd = 180 – 2P = 180 – 2(65) = 180 – 130 = 50

Q = 50

Equilibrium quantity = 50

 

a.). If the price is set at $72, it means it is above the equilibrium price of $65 and hence a price floor.

When a price floor is set above the equilibrium price, quantity supplied will exceed quantity demanded, and excess supply or surpluses will result.


The government should set a production quota below the equilibrium quantity at around 30 units to limit production and oversupply in the market.